# A galactic outflow traced by its extended Mg II emission out to a $\sim 30$ kpc radius in the Hubble Ultra Deep Field with MUSE

### v1

@@ 5.2 Modeling integral field spectroscopic data
Our goal is to produce a model cube that matches the data, for a given set of parameters. Thus, we compute the modeled absorption and emission profiles independently for each discrete sightline (i.e., MUSE spaxel), according to their respective impact parameter to the central source (also taking into account instrumental effects, see Sec. 5.5).
For a given MUSE spaxel, the fraction of the continuum flux absorbed at each wavelength by the spherical envelope located at resonance with the photons emitted by the central source at wavelength $\lambda$ is given by:
$I_{\rm abs}(\lambda)=1-e^{-\tau(\lambda)}$ (8)
Note that for a high optical depth, $I_{\rm abs}\approx 1$, meaning that the continuum emission is completely absorbed.
Resonant photons absorbed by the spherical envelope can then be re-emitted towards the observer. We calculate the emission produced by the gas in the envelope by assuming that the re-emitted photons are able to escape their resonant shell without further interactions (i.e., single scattering approximation). The profile of the emission for any given sightline, as a function of wavelength, can be then expressed in units of the continuum radiation as:
$I_{\rm ems}(\lambda)=\frac{1-e^{-\tau(\lambda)}}{4\pi r_{\lambda}^{2}}$ (9)
@@
However, note that while the absorbing component is scaled by the level of continuum in any specific sightline, the emission contribution must be scaled to the total continuum radiation emitted by the central source. This is because while the absorption is only meaningful relative to a continuum, the emission component does not depend on the presence of a continuum in the same sightline, but it is powered by the total production of continuum photons by the central source.
By construction, the absorption component is confined to wavelengths that correspond to $v_{\rm obs}<0$, since only the part of the spherical envelope moving towards the observer, between the observer and the central source, is able to absorb continuum radiation in the line of sight. On the other hand, both the blueshifted and redshifted parts of any given spherical shell are able to scatter photons into the line of sight. However, note that since $R_{0}$ represents the launching radius of the wind, wavelengths that correspond to a radius smaller than $R_{0}$ do not absorb or emit scattered radiation.
The resulting P-Cygni profile of the outflow can be computed for any specific sightline as:
$I(\lambda)=I_{\rm cont,local}(1-I_{\rm abs})+I_{\rm cont,total}\;I_{\rm ems}$ (10)
where $I_{\rm cont,local}$ and $I_{\rm cont,total}$ correspond to the local level of continuum in the sightline, and the total continuum of the galaxy in the corresponding wavelengths, respectively, with $R_{0}$, $\tau_{0}$, $\gamma$, $v_{0}$, and $v_{\rm max}$ being free parameters of the model. As explained in Sec. 4.1, the level of the continuum in each spaxel is calculated with a running median in the spectral direction of the data cube, within a window of $100$ pixels. This continuum is then passed to the model cube. Thus, by construction, the continuum in the data cube and model cube are identical.
@@
To avoid including unphysical Mg II absorption far away from the central galaxy, where the local level of continuum might be different from zero, either due to noise effects or due to scattered light from nearby (in projected distance) sources, we have set a maximum radius for absorption, beyond which the absorption is set to zero, even if the continuum is different from zero. This maximum absorption radius has been empirically determined as the radius where the radial profile computed in the absorption-dominated band of the P-Cygni profile (see Sec. 7.3 for details on how the absorption-dominated band is defined) intersects the continuum. That is, the radius beyond which there is no more Mg II absorption in the data cube. We find this radius to be $1\aas@@fstack{\prime\prime}1$, which corresponds to a physical distance of $\sim 8.2$ kpc at $z=0.737$.
@@ 5.3 Nebular emission contribution
Thus, we model this possible extra source of Mg II photons in a given sightline as a Gaussian (centered at 00 km s-1) whose amplitude is proportional to the local stellar continuum, through a free parameter $f_{C}$. Specifically, we define:
$I_{\rm Neb}(\lambda)=(f_{C}\times I_{\rm cont,local})e^{-\frac{1}{2}\frac{(\lambda-\lambda_{0})^{2}}{\sigma_{\rm Neb}^{2}}}$ (11)
@@
This additional component does not take into account possible in-place production of Mg II photons in the halo. This possibility is described in detail in Zabl et al. (2021). Although as described earlier in Sec. 4.1, the emission line ratios in the spectrum of #884 suggest that the extended Mg II emission primarily originates from the resonant scattering of continuum photons, we can not rule out a partial contribution of non-scattering photons (e.g., shocks), especially given the presence of extended emission from optical non-resonant lines (Wisotzki et al. in prep.).
@@ 5.4 Biconical geometry
The position angle corresponds to the angle of the outflow with respect to the horizontal axis, in the plane of the sky (i.e., perpendicular to the line of sight, see the left panel of Fig. 7). It is defined within the range $0\leq\mathrm{P.A.}\leq\pi$, where a P.A. of zero represents a completely horizontal outflow in the plane of the sky, and a P.A. of $\frac{\pi}{2}$ corresponds to the vertical case.
Finally, the rotation angle measures the rotation of the outflow in the plane defined by $x=0$, with respect to the $y$-axis. That is, a rotation perpendicular to the plane of the sky, towards the observer (see the right panel of Fig. 7). It is defined in the range $-\frac{\pi}{2}\leq\mathrm{R.A.}\leq\frac{\pi}{2}$, where zero represents an outflow in the plane of the sky, perpendicular to the line of sight, and both $-\frac{\pi}{2}$ and $\frac{\pi}{2}$ correspond to the case where the outflow is pointing directly to the observer.
Hence, for a given set of O.A., P.A. and R.A., we set the scattered emission outside the biconical volume to zero, whereas inside the outflow it is given by Eq. 9
Figure 7: Schematic representation of the outflow geometry parametrization described in Sec. 5.4. The figure illustrates the front (left) and side (right) view of an outflow characterized by an arbitrary O.A., P.A. $=90^{\circ}$ and R.A. $=45^{\circ}$. The opening angle represents the angle between the axis and the surface of each outflow cone, the position angle corresponds to the angle of the outflow with respect to the horizontal axis, in the plane of the sky, and the rotation angle measures the rotation of the outflow in the plane defined by $x=0$, with respect to the $y$-axis. The red line shows the cone axis. The observer is located at a large positive $z$ value.
5.5 PSF and LSF convolution
Accounting for the instrumental effects on the spectral line profile and on the spatial light distribution is essential for a meaningful comparison between our model, described in the previous subsections, and the data. These nonphysical effects are encoded by the instrumental line spread function (LSF) and the point spread function (PSF), where the first describes the spectral broadening of a given spectral line due to instrumental effects, and the second characterizes the spatial broadening of the light distribution of a point like source in the detector. Both the PSF and LSF of the MUSE-HUDF are well characterized in Bacon et al. (2017), and we use the same parametrization to model these effects. Specifically, we apply an equivalent spectral and spatial convolution to the modeled datacube before comparing it with the observed datacube. Here, we provide a brief summary of our approach to account for these effects, but we refer the reader to Bacon et al. (2017) for a more detailed description.
@@ 5.5.1 Line spread function
The LSF is modeled as a Gaussian function, whose width was measured by fitting a Gaussian to a set of sky lines, accros all spaxels in the mosaic. Bacon et al. (2017) find that the FWHM of the LSF varies smoothly with wavelength, as a second degree polynomial:
$\mathrm{FWHM}(\lambda)=5.835\times 10^{-8}\lambda^{2}-9.080\times 10^{-4}\lambda+5.983$ (12)
@@ 5.5.2 Point spread function
During commissioning, a detailed analysis of the MUSE PSF showed that it was very well modeled by a Moffat circular function with constant $\beta$ and a FWHM that varies linearly with wavelength. The values of the FWHM and $\beta$ were determined by matching the resolution of a set HST broadband images, and analog synthetic images constructed from the MUSE datacubes, using the same spectral response and total flux calibration as for the HST images, provided that the HST PSF is well known [^3] This methodology was adopted because most of MUSE fields do not have bright stars to model the PSF independently.. Bacon et al. (2017) find that the PSF is well represented by a Moffat with $\beta=2.8$ and FWHM that decreases with wavelength, from $0\aas@@fstack{\prime\prime}71$ at the blue end, to $0\aas@@fstack{\prime\prime}57$ at the red end. The observed wavelength of the Mg II transition is very close to the blue end of the MUSE wavelength range ($\sim 4860$ $\AA$), which corresponds to a FWHM of $\sim 0\aas@@fstack{\prime\prime}7$ for the PSF.
@@ 6.1 Modeled signal
We use the model described in the previous section to reproduce the Mg II P-Cygni profile exhibited by the galaxy MUSE-HUDF #$\#884$. To do this, we consider the wavelength interval of 2790 - 2809 $\AA$ (rest-frame), thus modeling both Mg II transitions simultaneously (see Fig 4). In order to avoid contamination from nearby galaxies (in projected distance, particularly from BG1 and FG2, see Sec. 3.1), we fit the Mg II emission profile in a circular field of view with a radius of $3\aas@@fstack{\prime\prime}5$ ($\sim$26 kpc) around the central galaxy (blue aperture in Fig 3).
@@ 6.2 Description of the fitting
We used the following priors for the outflow parameters, described in the Sec. 5.1:
@@
Note that for $R_{0}$ we chose a normal prior rather than a Uniform one. This is because we know the approximate size of the galaxy from HST photometry (see Sec. 1), and this value should correspond approximately to $R_{0}$ (only approximately because $R_{0}$ represents a simplification of the model, assuming that the continuum central source is spherical). Thus, we exclude scenarios where $R_{0}$ is much larger or much smaller than the actual size of #884. Furthermore, given the MUSE spatial and spectral resolution, $R_{0}$ is only partially constrained, so we purposely use narrow priors around the half-light radius measurement, instead of using a fixed $R_{0}$ value.
We have also included an extra parameter $\Delta v$ to account for any kinematic offset between the systemic velocity and the P-Cygni emission. Its prior is given by:
$\Delta v\sim\mathcal{U}(80,130)\$ (19)
and it has units of km s-1.
For the geometric parameters, described in Sec. 5.4, we considered the following priors:
@@
in units of degrees. Note that for P.A. we also chose a Normal prior, instead of a more agnostic Uniform distribution. This choice is justified on the basis that the Mg II pseudo-narrow band image extracted from the MUSE datacube (see Fig. 3) shows a preferred direction on the spatial distribution of the Mg II emission, along the diagonal axis, from the top-left to the bottom-right corner of the image.
Finally, we included an additional term $f$ to account for the existence of additional variance in the data. The extra variance ($\sigma$) is proportional to the model, such that:
$\sigma^{2}=f^{2}M^{2}$ (23)
where $M$ is the value of the model, evaluated in the independent variable, given a set of parameters. The prior on $f$ is given by:
$f\sim\mathcal{L}\mathcal{U}(-5,-1)$ (24)
@@ 7.1 Best-fitting model
In this section, we investigate the best-fitting model obtained using the fitting routine described in the previous section. Figure 8 shows a pseudo-narrow band image convolved with a Gaussian of 1″ FWHM, for the MUSE data cube and the best-fitting model cube, collapsing the Mg II wavelength range, for the field of view where we compute the model (in the case of the model cube, this convolution is applied in addition to the MUSE PSF, such that both cubes are similarly convolved). The stellar continuum has been subtracted from both images to better compare the morphology of the Mg II absorption and emission components.
@@
Figure 9: Posterior distributions of each one of the model parameters described in Sec. 5, and their respective covariance. The median value of each posterior distribution is indicated in the top panels. The uncertainties show the percentiles 32 and 68 of the distribution.
Figure 9 shows the posterior distributions of each of the model parameters described in Sec. 5, and their respective covariance.
Despite the inability of MUSE to spatially resolve $R_{0}$, the posterior distribution of $R_{0}$ is not centered around the prior value of 1.5 kpc. Instead, the posterior distribution is shifted towards a slightly larger radii ($R_{0}=1.6\pm 0.2$ kpc), showing that a continuum source size larger than 1.5 kpc is marginally favored to reproduce the data (although 1.5 kpc is within the error bars of the best-fitting value). However, we acknowledge that the width of this posterior distribution is partially driven by the narrow prior selection and that $R_{0}$ is only marginally constrained in our dataset.
On the other hand, the posterior distribution of the optical depth $\tau_{0}$ exhibits a clear maximum at $\tau_{0}=4000\pm 2000$ (albeit with a tail towards even higher values, hence, the large uncertainty in the parameter). It is remarkable that despite the heavily saturated nature of the Mg II absorption line, our modeling scheme still allows us to put a constraint on the gas optical depth. This is possible because although the absorption is completely saturated in the inner part of the outflow, our model is still sensitive to the changes in the optical depth towards larger radii, that is, the radial gradient of $\tau$. As a consequence, we find some degree of covariance between $\tau_{0}$ and the quantities that determine the velocity (and $\tau$) gradient, $\gamma$ and $v_{0}$. The combination of parameters of our best-fitting model leads to $\tau=1$ only at $R\sim 60$ kpc. In principle, beyond this radius, we would not need the Sobolev approximation since most photons would simply escape the galaxy without scattering with the ions in the CGM, independently of the velocity gradient. However, this only occurs at radii much larger than the impact parameters considered in our model, and therefore it does not represent a conflict with our assumptions.
@@
However, a wind velocity that increases with radius does not necessarily imply that the wind is indeed accelerating. If there is a distribution of velocities for the gas leaving the ISM, then higher velocity material will naturally dominate at larger radii, even if the wind is decelerating at all radii, as shown by the hydrodynamical simulations of SNe feedback-powered outflows from Dalla Vecchia & Schaye (2008), producing a velocity profile that increases with radii. We caution the reader about this possible degeneracy in order to avoid an overinterpretation of our modeling results.
The P-Cygni profile shows a velocity offset $\Delta v=97\pm 5$ km s-1 with respect to the systemic redshift. This offset can be appreciated in Fig. 4, especially in the innermost apertures, where the Mg II rest-frame wavelength coincides with the absorption, rather than with the emission peak. On the other hand, in the outer apertures, the Mg II rest-frame wavelength roughly coincides with the emission line. Thus, this inferred velocity offset is likely a result of the complex kinematic structure of the outflow not being properly described by a single power law. We discuss this further in Sec. 8.
@@
Regarding the nebular contribution to the emission, parameterized by $f_{C}$, we find that a minor contribution is needed to model the data ($\sim 12\%$ of the continuum level). However, this might not be entirely correct, since our model is not able to reproduce the observed ratio of the Mg II emission lines, particularly in the inner apertures, where nebular emission could be present. This limitation is further discussed in Sec. 8.
Finally, the geometry of the outflow is well constrained by our model. We find a nearly face-on outflow (R.A. = $80\pm 10$), meaning that we observe the cone almost along its symmetry axis ‘down the barrel’. This coincides with what we would expect from a polar outflow, since #884 is also seen nearly face-on. Therefore, P.A. = $120\pm 10$ has little impact on the fitting results. The resulting O.A. of the bicone is large, $59\pm 4$, and this produces an asymmetric emission line profile that is particularly evident in the outermost apertures in Fig. 11. In general, the inferred geometry resembles what we see in the data cube, as shown in Fig. 8. Figure 10 provides a schematic representation of the biconical geometry produced by this combination of angles. Table 2 shows the mean and standard deviation of the posterior distribution of all our model parameters.
Table 2: Summary of best-fitting model parameters
@@
Figure 10: Schematic representation of the geometry of our best-fitting outflow model, with an O.A. = 59, P.A. = 120, and R.A. = 80, as described in Sec. 7.1. The red dashed line indicates the cone axis. The observer is located at a large positive $z$ value.

### v3

@@ 5.2 Modeling integral field spectroscopic data
Our goal is to produce a model cube that matches the data, for a given set of parameters. Thus, we compute the modeled absorption and emission profiles independently for each discrete sight line (i.e., MUSE spaxel), according to their respective impact parameter to the central source (also taking into account instrumental effects; see Sect. 5.5).
For a given MUSE spaxel, the fraction of the continuum flux absorbed at each wavelength by the spherical envelope located at resonance with the photons emitted by the central source at wavelength $\lambda$ is given by
$I_{\rm abs}(\lambda)=1-e^{-\tau(\lambda)}.$ (8)
We note that for a high optical depth, $I_{\rm abs}\approx 1$, meaning that the continuum emission is completely absorbed.
Resonant photons absorbed by the spherical envelope can then be re-emitted toward the observer. We calculate the emission produced by the gas in the envelope by assuming that the re-emitted photons are able to escape their resonant shell without further interactions (i.e., single scattering approximation). The profile of the emission for any given sight line, as a function of wavelength, can be then expressed in units of the continuum radiation as
$I_{\rm ems}(\lambda)=\frac{1-e^{-\tau(\lambda)}}{4\pi r_{\lambda}^{2}},$ (9)
@@
However, we note that while the absorbing component is scaled by the level of continuum in any specific sight line, the emission contribution must be scaled to the total continuum radiation emitted by the central source. This is because while the absorption is only meaningful relative to a continuum, the emission component does not depend on the presence of a continuum in the same sight line, but it is powered by the total production of continuum photons by the central source.
By construction, the absorption component is confined to wavelengths that correspond to $v_{\rm obs}<0$, since only the part of the spherical envelope moving toward the observer, between the observer and the central source, is able to absorb continuum radiation in the line of sight. On the other hand, both the blueshifted and redshifted parts of any given spherical shell are able to scatter photons into the line of sight. However, we note that since $R_{0}$ represents the launching radius of the wind, wavelengths that correspond to a radius smaller than $R_{0}$ do not absorb or emit scattered radiation.
The resulting P-Cygni profile of the outflow can be computed for any specific sight line as
$I(\lambda)=I_{\rm cont,local}(1-I_{\rm abs})+I_{\rm cont,total}\;I_{\rm ems},$ (10)
where $I_{\rm cont,local}$ and $I_{\rm cont,total}$ correspond to the local level of continuum in the sight line, and the total continuum of the galaxy in the corresponding wavelengths, respectively, with $R_{0}$, $\tau_{0}$, $\gamma$, $v_{0}$, and $v_{\rm max}$ being free parameters of the model. As explained in Sect. 4.1, the level of the continuum in each spaxel is calculated with a running median in the spectral direction of the data cube, within a window of $100$ pixels. This continuum is then passed to the model cube. Thus, by construction, the continuum in the data cube and model cube are identical.
@@
To avoid including unphysical Mg II absorption far away from the central galaxy, where the local level of continuum might be different from zero, either due to noise effects or due to scattered light from nearby (in projected distance) sources, we have set a maximum radius for absorption, beyond which the absorption is set to zero, even if the continuum is different from zero. This maximum absorption radius has been empirically determined as the radius where the radial profile computed in the absorption-dominated band of the P-Cygni profile (see Sect. 7.3 for details on how the absorption-dominated band is defined) intersects the continuum. That is, the radius beyond which there is no more Mg II absorption in the data cube. We find this radius to be $1\aas@@fstack{\prime\prime}1$, which corresponds to a physical distance of $\sim 8.2$ kpc at $z=0.737$.
@@ 5.3 Nebular emission contribution
Thus, we modeled this possible extra source of Mg II photons in a given sight line as a Gaussian (centered at 00 km s-1) whose amplitude is proportional to the local stellar continuum, through a free parameter $f_{C}$. Specifically, we define
$I_{\rm Neb}(\lambda)=(f_{C}\times I_{\rm cont,local})e^{-\frac{1}{2}\frac{(\lambda-\lambda_{0})^{2}}{\sigma_{\rm Neb}^{2}}},$ (11)
@@
This additional component does not take into account possible in-place production of Mg II photons in the halo. This possibility is described in detail in Zabl et al. (2021). Although as described earlier in Sect. 4.1, the emission line ratios in the spectrum of #884 suggest that the extended Mg II emission primarily originates from the resonant scattering of continuum photons, we cannot rule out a partial contribution of non-scattering photons (e.g., shocks), especially given the presence of extended emission from optical non-resonant lines (Wisotzki et al. in prep.).
@@ 5.4 Biconical geometry
The position angle corresponds to the angle of the outflow with respect to the horizontal axis, in the plane of the sky (i.e., perpendicular to the line of sight; see the left panel of Fig. 7). It is defined within the range $0\leq\mathrm{P.A.}\leq\pi$, where a P.A. of zero represents a completely horizontal outflow in the plane of the sky, and a P.A. of $\frac{\pi}{2}$ corresponds to the vertical case.
Finally, the rotation angle measures the rotation of the outflow in the plane defined by $x=0$, with respect to the $y$-axis. That is, a rotation perpendicular to the plane of the sky, toward the observer (see the right panel of Fig. 7). It is defined in the range $-\frac{\pi}{2}\leq\mathrm{R.A.}\leq\frac{\pi}{2}$, where zero represents an outflow in the plane of the sky, perpendicular to the line of sight, and both $-\frac{\pi}{2}$ and $\frac{\pi}{2}$ correspond to the case where the outflow is pointing directly to the observer.
Hence, for a given set of O.A., P.A., and R.A., we set the scattered emission outside the biconical volume to zero, whereas inside the outflow it is given by Eq. 9
Figure 7: Schematic representation of the outflow geometry parametrization described in Sect. 5.4. The figure illustrates the front (left) and side (right) view of an outflow characterized by an arbitrary O.A., P.A. $=90^{\circ}$ and R.A. $=45^{\circ}$. The opening angle represents the angle between the axis and the surface of each outflow cone, the position angle corresponds to the angle of the outflow with respect to the horizontal axis, in the plane of the sky, and the rotation angle measures the rotation of the outflow in the plane defined by $x=0$, with respect to the $y$-axis. The red line shows the cone axis. The observer is located at a large positive $z$ value.
5.5 Line spread function and point spread function convolution
Accounting for the instrumental effects on the spectral line profile and on the spatial light distribution is essential for a meaningful comparison between our model, described in the previous subsections, and the data. These nonphysical effects are encoded by the instrumental LSF and PSF, where the first describes the spectral broadening of a given spectral line due to instrumental effects, and the second characterizes the spatial broadening of the light distribution of a point-like source in the detector. Both the LSF and PSF of the MUSE-HUDF are well characterized in Bacon et al. (2017), and we use the same parametrization to model these effects. Specifically, we apply an equivalent spectral and spatial convolution to the modeled data cube before comparing it with the observed data cube. Here, we provide a brief summary of our approach to account for these effects, but we refer the reader to Bacon et al. (2017) for a more detailed description.
@@ 5.5.1 Line spread function
The LSF is modeled as a Gaussian function, whose width was measured by fitting a Gaussian to a set of sky lines, across all spaxels in the mosaic. Bacon et al. (2017) find that the FWHM of the LSF varies smoothly with wavelength, as a second-degree polynomial:
$\mathrm{FWHM}(\lambda)=5.835\times 10^{-8}\lambda^{2}-9.080\times 10^{-4}\lambda+5.983,$ (12)
@@ 5.5.2 Point spread function
During commissioning, a detailed analysis of the MUSE PSF showed that it was very well modeled by a Moffat circular function with constant $\beta$ and a FWHM that varies linearly with wavelength. The values of the FWHM and $\beta$ were determined by matching the resolution of a set of HST broadband images, and analog synthetic images constructed from the MUSE data cubes, using the same spectral response and total flux calibration as for the HST images, provided that the HST PSF is well known [^3] This methodology was adopted because most of MUSE fields do not have bright stars to model the PSF independently.. Bacon et al. (2017) find that the PSF is well represented by a Moffat with $\beta=2.8$ and FWHM that decreases with wavelength, from $0\aas@@fstack{\prime\prime}71$ at the blue end, to $0\aas@@fstack{\prime\prime}57$ at the red end. The observed wavelength of the Mg II transition is very close to the blue end of the MUSE wavelength range ($\sim 4860$ $\AA$), which corresponds to a FWHM of $\sim 0\aas@@fstack{\prime\prime}7$ for the PSF.
@@ 6.1 Modeled signal
We use the model described in the previous section to reproduce the Mg II P-Cygni profile exhibited by the galaxy MUSE-HUDF #$\#884$. To do this, we consider the wavelength interval of 2790 - 2809 $\AA$ (rest-frame), thus modeling both Mg II transitions simultaneously (see Fig 4). In order to avoid contamination from nearby galaxies (in projected distance, particularly from BG1 and FG2; see Sect. 3.1), we fit the Mg II emission profile in a circular FoV with a radius of $3\aas@@fstack{\prime\prime}5$ ($\sim$26 kpc) around the central galaxy (blue aperture in Fig 3).
@@ 6.2 Description of the fitting
We used the following priors for the outflow parameters, described in Sect. 5.1:
@@
We note that for $R_{0}$ we chose a normal prior rather than a uniform one. This is because we know the approximate size of the galaxy from HST photometry (see Sect. 1), and this value should correspond approximately to $R_{0}$ (only approximately because $R_{0}$ represents a simplification of the model, assuming that the continuum central source is spherical). Thus, we exclude scenarios where $R_{0}$ is much larger or much smaller than the actual size of #884. Furthermore, given the MUSE spatial and spectral resolution, $R_{0}$ is only partially constrained, so we purposely use narrow priors around the half-light radius measurement, instead of using a fixed $R_{0}$ value.
We have also included an extra parameter $\Delta v$ to account for any kinematic offset between the systemic velocity and the P-Cygni emission. Its prior is given by
$\Delta v\sim\mathcal{U}(80,130)\\ ,$ (19)
and it has units of kilometers per second.
For the geometric parameters, described in Sect. 5.4, we considered the following priors:
@@
in units of degrees. We note that for P.A. we also chose a normal prior, instead of a more agnostic uniform distribution. This choice is justified on the basis that the Mg II pseudo-narrowband image extracted from the MUSE data cube (see Fig. 3) shows a preferred direction on the spatial distribution of the Mg II emission, along the diagonal axis, from the top-left to the bottom-right corner of the image.
Finally, we included an additional term $f$ to account for the existence of additional variance in the data. The extra variance ($\sigma$) is proportional to the model such that
$\sigma^{2}=f^{2}M^{2},$ (23)
where $M$ is the value of the model, evaluated in the independent variable, given a set of parameters. The prior on $f$ is given by
$f\sim\mathcal{L}\mathcal{U}(-5,-1).$ (24)
@@ 7.1 Best-fitting model
In this section, we investigate the best-fitting model obtained using the fitting routine described in the previous section. Figure 8 shows a pseudo-narrowband image convolved with a Gaussian of 1″ FWHM, for the MUSE data cube and the best-fitting model cube, collapsing the Mg II wavelength range, for the FoV where we compute the model (in the case of the model cube, this convolution is applied in addition to the MUSE PSF such that both cubes are similarly convolved). The stellar continuum has been subtracted from both images to better compare the morphology of the Mg II absorption and emission components.
@@
Figure 9: Posterior distributions of each one of the model parameters described in Sect. 5 and their respective covariance. The median value of each posterior distribution is indicated in the top panels. The uncertainties show the percentiles 32 and 68 of the distribution.
Figure 9 shows the posterior distributions of each of the model parameters described in Sect. 5, and their respective covariance.
Despite the inability of MUSE to spatially resolve $R_{0}$, the posterior distribution of $R_{0}$ is not centered around the prior value of 1.5 kpc. Instead, the posterior distribution is shifted toward a slightly larger radii ($R_{0}=1.6\pm 0.2$ kpc), showing that a continuum source size larger than 1.5 kpc is marginally favored to reproduce the data (although 1.5 kpc is within the error bars of the best-fitting value). However, we acknowledge that the width of this posterior distribution is partially driven by the narrow prior selection and that $R_{0}$ is only marginally constrained in our dataset.
On the other hand, the posterior distribution of the optical depth $\tau_{0}$ exhibits a clear maximum at $\tau_{0}=4000\pm 2000$ (albeit with a tail toward even higher values, hence, the large uncertainty in the parameter). It is remarkable that despite the heavily saturated nature of the Mg II absorption line, our modeling scheme still allows us to put a constraint on the gas optical depth. This is possible because although the absorption is completely saturated in the inner part of the outflow, our model is still sensitive to the changes in the optical depth toward larger radii, that is, the radial gradient of $\tau$. As a consequence, we find some degree of covariance between $\tau_{0}$ and the quantities that determine the velocity (and $\tau$) gradient, $\gamma$ and $v_{0}$. The combination of parameters of our best-fitting model leads to $\tau=1$ only at $R\sim 60$ kpc. In principle, beyond this radius, we would not need the Sobolev approximation since most photons would simply escape the galaxy without scattering with the ions in the CGM, independently of the velocity gradient. However, this only occurs at radii much larger than the impact parameters considered in our model, and therefore, it does not represent a conflict with our assumptions.
@@
However, a wind velocity that increases with radius does not necessarily imply that the wind is indeed accelerating. If there is a distribution of velocities for the gas leaving the interstellar medium (ISM), then higher velocity material will naturally dominate at larger radii, even if the wind is decelerating at all radii, as shown by the hydrodynamical simulations of SNe feedback-powered outflows from Dalla Vecchia & Schaye (2008), producing a velocity profile that increases with radii. We caution the reader about this possible degeneracy in order to avoid an overinterpretation of our modeling results.
The P-Cygni profile shows a velocity offset $\Delta v=97\pm 5$ km s-1 with respect to the systemic redshift. This offset can be appreciated in Fig. 4, especially in the innermost apertures, where the Mg II rest-frame wavelength coincides with the absorption, rather than with the emission peak. On the other hand, in the outer apertures, the Mg II rest-frame wavelength roughly coincides with the emission line. Thus, this inferred velocity offset is likely a result of the complex kinematic structure of the outflow not being properly described by a single power law. We discuss this further in Sect. 8.
@@
Regarding the nebular contribution to the emission, parameterized by $f_{C}$, we find that a minor contribution is needed to model the data ($\sim 12\%$ of the continuum level). However, this might not be entirely correct, since our model is not able to reproduce the observed ratio of the Mg II emission lines, particularly in the inner apertures, where nebular emission could be present. This limitation is further discussed in Sect. 8.
Finally, the geometry of the outflow is well constrained by our model. We find a nearly face-on outflow (R.A. = $80\pm 10$), meaning that we observe the cone almost along its symmetry axis ”down the barrel”. This coincides with what we would expect from a polar outflow, since #884 is also seen nearly face-on. Therefore, P.A. = $120\pm 10$ has little impact on the fitting results. The resulting O.A. of the bicone is large, $59\pm 4$, and this produces an asymmetric emission line profile that is particularly evident in the outermost apertures in Fig. 11. In general, the inferred geometry resembles what we see in the data cube, as shown in Fig. 8. Figure 10 provides a schematic representation of the biconical geometry produced by this combination of angles. Table 2 shows the mean and standard deviation of the posterior distribution of all our model parameters.
Table 2: Best-fitting model parameters.
@@
Figure 10: Schematic representation of the geometry of our best-fitting outflow model, with O.A. = 59, P.A. = 120, and R.A. = 80, as described in Sect. 7.1. The red dashed line indicates the cone axis. The observer is located at a large positive $z$ value.